Hello.
I'm needing some help with checking what I did wrong on this problem:
Factor each completely. Some may require several steps
3n^2+14n-5
thanks!
n=[-14+ root((14^2 - 4*3*(-5))]/2*3 = [-14 + root(196 + 60)]/6 = [-14+16]/6= 1/3
n=[-14- root((14^2 - 4*3*(-5))]/2*3 = [-14 - root(196 + 60)]/6 = [-14-16]/6= -5
The result is n=1/3 and n=-5

Found 2 solutions by Earlsdon, dukesmart:
Answer by Earlsdon(6294)
Answer by dukesmart(20)

=[-14+ sqrt((14^2 - 4*3*(-5))]/2*3 = [-14 + root(196 + 60)]/6 = [-14+16]/6= 1/3
n=[-14- sqrt((14^2 - 4*3*(-5))]/2*3 = [-14 - root(196 + 60)]/6 = [-14-16]/6= -5
The result is n=1/3 and n=-5
so I can write:
3n^2+14n-5=
= (3n-1) (n+5)
I solve that quadratic equation in case you encounter any complicated problem , you can simplify it. ((any)